# Adaptive-TMLE for the Average Treatment Effect based on Randomized Controlled Trial Augmented with Real-World Data

### v1

@@ Acknowledgement
Let ${\cal M}_{w,n}$ be a fixed working model (learned on external data set). Let $\psi_{n}^{*}=1/V\sum_{v=1}^{V}\Psi_{{\cal M}_{w,n}}(P_{n,v}^{*})$ be the CV-TMLE of $1/V\sum_{v=1}^{V}\Psi_{{\cal M}_{w,n}}(P_{0})$ satisfying $1/V\sum_{v}P_{n,v}^{1}D_{{\cal M}_{w,n},P_{n,v}^{*}}=o_{P}(n^{-1/2})$. We have
$\psi_{n}^{*}-\Psi_{{\cal M}_{w,n}}(P_{0})=\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n},P_{n,v}^{*}}+\frac{1}{V}\sum_{v=1}^{V}R_{{\cal M}_{w,n}}(P_{n,v}^{*},P_{0})+o_{P}(n^{-1/2}).$
@@
$\Psi_{{\cal M}_{w,n}}(P_{0})-\Psi(P_{0})=o_{P}(n^{-1/2}).$
A sufficient condition for this is presented in Theorem 1: Let $P_{0,n}\equiv\Pi_{{\cal M}_{w,n}}(P_{0})$; $\tilde{P}_{0,n}=\Pi(P_{0,n}\mid{\cal M}_{0})$ for a submodel ${\cal M}_{0}\subset{\cal M}$ containing $P_{0}$. Let $\tilde{D}_{{\cal M}_{w,n},P_{0,n}}$ be the projection of $D_{{\cal M}_{0},P_{0,n}}$ onto the tangent space $T_{{\cal M}_{w,n}}(P_{0,n})\subset L^{2}_{0}(P_{0,n})$ of the working model ${\cal M}_{w,n}$ at $P_{0,n}$. Assume that ${\cal M}_{w,n}$ approximates $P_{0}$ in the sense that $\Psi(\tilde{P}_{0,n})-\Psi(P_{0,n})=o_{P}(n^{-1/2})$ and
$(\tilde{P}_{0,n}-P_{0})\left\{D_{{\cal M}_{0},\tilde{P}_{0,n}}-\tilde{D}_{{\cal M}_{w,n},\tilde{P}_{0,n}}\right\}=o_{P}(n^{-1/2}).$
@@
$\psi_{n}^{*}-\Psi(P_{0})=\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w,n},P_{n,v}^{*}}+\frac{1}{V}\sum_{v=1}^{V}R_{{\cal M}_{w,n}}(P_{n,v}^{*},P_{0})+o_{P}(n^{-1/2}).$
Rate of convergence condition: Assume $1/V\sum_{v=1}^{V}R_{{\cal M}_{w,n}}(P_{n,v}^{*},P_{0})=o_{P}(n^{-1/2})$.
Weak Donsker class condition: Assume, conditional on the training sample $P_{n,v}$, $D_{{\cal M}_{w,n,v},P_{n,v}^{*}}$ falls in a $P_{0}$-Donsker class (note that only the targeting step depends on $P_{n,v}^{1}$, making this a very weak condition). Then,
$\psi_{n}^{*}-\Psi(P_{0})=O_{P}(n^{-1/2}).$
Consistency of TMLE $P_{n,v}^{*}$ to $P_{0}$: Suppose $\max_{v=1}^{V}\parallel D_{{\cal M}_{w,n},P_{n,v}^{*}}-D_{{\cal M}_{w,n},P_{0}}\parallel_{P_{0}}=o_{P}(1)$. Then,
$\psi_{n}^{*}-\Psi(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w,n},P_{0}}+o_{P}(n^{-1/2}).$
@@
$\sigma_{w,n}^{-1}(\psi_{n}^{*}-\psi_{0})\Rightarrow_{d}N(0,1),$
where $\sigma_{w,n}^{2}=P_{0}D_{{\cal M}_{w,n},P_{0}}^{2}$.
Asymptotic linearity condition: Assume $\parallel D_{{\cal M}_{w,n},P_{0}}-D_{{\cal M}_{0},P_{0}}\parallel_{P_{0}}=o_{P}(1)$. Under this stronger condition on ${\cal M}_{w,n}$, we have asymptotic linearity with a super-efficient influence curve:
$\psi_{n}^{*}-\Psi(P_{0})=P_{n}D_{{\cal M}_{0},P_{0}}+o_{P}(n^{-1/2}).$
@@
To demonstrate this proposal we consider a statistical model that corresponds with variation independent models of conditional densities. Specifically, let $O=(O_{1},\ldots,O_{J})$ and consider a statistical model ${\cal M}$ that assumes $p(O)=\prod_{j=1}^{J}p_{O(j)\mid Pa(O_{j})}(O(j)\mid Pa(O_{j}))$, but leaves all the conditional densities of $O(j)$, given its parents $Pa(O_{j})$, unspecified, $j=1,\ldots,J$. For example, the likelihood might be factorized according to the ordering $p(O)=\prod_{j=1}^{J}p(O(j)\mid\bar{O}(j-1))$, where $\bar{O}(j-1)=(O(1),\ldots,O(j-1))$, and one might not make any assumptions, or one might assume that for some of the conditional densities $p(O(j)\mid\bar{O}(j-1))$ it is known to depend on $\bar{O}(j-1)$ through a dimension reduction $Pa(O(j))$. This defines then a statistical model ${\cal M}$ for $P_{0}$ only driven by conditional independence assumptions.
Working model defined by estimated dimension reductions: For a given $j$, let $S_{n,j}=S_{j}(P_{n})$ be a data dependent dimension reduction in the sense that $S_{n,j}(Pa(O_{j}))$ is of (much) smaller dimension than $Pa(O_{j})$. Consider the submodel ${\cal M}_{w,n}$ defined by assuming $p_{O(j)\mid Pa(O_{j})}=p_{O(j)\mid S_{n,j}(Pa(O_{j}))}$, $j=1,\ldots,J$: in other words, this working model assumes that $O(j)$ is independent of $Pa(O_{j})$, given the reduction $S_{n,j}(Pa(O_{j}))$, $j=1,\ldots,J$. Clearly, we have ${\cal M}_{w,n}\subset{\cal M}$ by making stronger conditional independence assumptions than the ones defining ${\cal M}$.
The Kullback-Leibler projection of $p$ onto ${\cal M}_{w,n}$ would involve projecting each $p_{j}=p_{O(j)\mid Pa(O)j)}\in{\cal M}_{j}$ onto its smaller working model ${\cal M}_{j,n}$ as follows:
@@
where $L(P)=-\log p$ is the log-likelihood loss. So this projection computes the MLE of $P_{j}$ over the working model under an infinite sample from $P$. Therefore, we have a clear definition of $\Pi_{P}(P\mid{\cal M}_{w,n})$ and also a clear understanding of how one estimates this projection $\Pi_{P_{0}}(P_{0}\mid{\cal M}_{w,n})$ with an MLE $\tilde{P}_{n}=\arg\min_{P_{1}\in{\cal M}_{w,n}}P_{n}L(P_{1})$ over ${\cal M}_{w,n}$, or, if ${\cal M}_{w,n}$ is too high dimensional, with a regularized MLE such as an HAL-MLE. If $Pa(O(j))$ is very high dimensional, then an HAL-estimator of the conditional density $p_{0,j}$ is cumbersome, while, if ${\cal M}_{w,n}$ is indexed by a low dimensional unspecified function, then we can estimate this projection with a powerful HAL-MLE that is also computationally very feasible.

### v2

@@ Acknowledgement
Let ${\cal M}_{w}$ be a fixed working model (learned on external data set). Let $\psi_{n}^{\star}=1/V\sum_{v=1}^{V}\Psi_{{\cal M}_{w}}(P_{n,v}^{\star})$ be the CV-TMLE of $1/V\sum_{v=1}^{V}\Psi_{{\cal M}_{w}}(P_{0})$ satisfying $1/V\sum_{v}P_{n,v}^{1}D_{{\cal M}_{w},P_{n,v}^{\star}}=o_{P}(n^{-1/2})$. We have
$\psi_{n}^{\star}-\Psi_{{\cal M}_{w}}(P_{0})=\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w},P_{n,v}^{\star}}+\frac{1}{V}\sum_{v=1}^{V}R_{{\cal M}_{w}}(P_{n,v}^{\star},P_{0})+o_{P}(n^{-1/2}).$
@@
$\Psi_{{\cal M}_{w}}(P_{0})-\Psi(P_{0})=o_{P}(n^{-1/2}).$
A sufficient condition for this is presented in Theorem 1: Let $P_{0,n}\equiv\Pi_{{\cal M}_{w}}(P_{0})$; $\tilde{P}_{0,n}=\Pi(P_{0,n}\mid{\cal M}_{0})$ for a submodel ${\cal M}_{0}\subset{\cal M}$ containing $P_{0}$. Let $\tilde{D}_{{\cal M}_{w},P_{0,n}}$ be the projection of $D_{{\cal M}_{0},P_{0,n}}$ onto the tangent space $T_{{\cal M}_{w}}(P_{0,n})\subset L^{2}_{0}(P_{0,n})$ of the working model ${\cal M}_{w}$ at $P_{0,n}$. Assume that ${\cal M}_{w}$ approximates $P_{0}$ in the sense that $\Psi(\tilde{P}_{0,n})-\Psi(P_{0,n})=o_{P}(n^{-1/2})$ and
$(\tilde{P}_{0,n}-P_{0})\left\{D_{{\cal M}_{0},\tilde{P}_{0,n}}-\tilde{D}_{{\cal M}_{w},\tilde{P}_{0,n}}\right\}=o_{P}(n^{-1/2}).$
@@
$\psi_{n}^{\star}-\Psi(P_{0})=\frac{1}{V}\sum_{v=1}^{V}(P_{n,v}^{1}-P_{0})D_{{\cal M}_{w},P_{n,v}^{\star}}+\frac{1}{V}\sum_{v=1}^{V}R_{{\cal M}_{w}}(P_{n,v}^{\star},P_{0})+o_{P}(n^{-1/2}).$
Rate of convergence condition: Assume $1/V\sum_{v=1}^{V}R_{{\cal M}_{w}}(P_{n,v}^{\star},P_{0})=o_{P}(n^{-1/2})$.
Weak Donsker class condition: Assume, conditional on the training sample $P_{n,v}$, $D_{{\cal M}_{w,n,v},P_{n,v}^{\star}}$ falls in a $P_{0}$-Donsker class (note that only the targeting step depends on $P_{n,v}^{1}$, making this a very weak condition). Then,
$\psi_{n}^{\star}-\Psi(P_{0})=O_{P}(n^{-1/2}).$
Consistency of TMLE $P_{n,v}^{\star}$ to $P_{0}$: Suppose $\max_{v=1}^{V}\parallel D_{{\cal M}_{w},P_{n,v}^{\star}}-D_{{\cal M}_{w},P_{0}}\parallel_{P_{0}}=o_{P}(1)$. Then,
$\psi_{n}^{\star}-\Psi(P_{0})=(P_{n}-P_{0})D_{{\cal M}_{w},P_{0}}+o_{P}(n^{-1/2}).$
@@
$\sigma_{w,n}^{-1}(\psi_{n}^{\star}-\psi_{0})\Rightarrow_{d}N(0,1),$
where $\sigma_{w,n}^{2}=P_{0}D_{{\cal M}_{w},P_{0}}^{2}$.
Asymptotic linearity condition: Assume $\parallel D_{{\cal M}_{w},P_{0}}-D_{{\cal M}_{0},P_{0}}\parallel_{P_{0}}=o_{P}(1)$. Under this stronger condition on ${\cal M}_{w}$, we have asymptotic linearity with a super-efficient influence curve:
$\psi_{n}^{\star}-\Psi(P_{0})=P_{n}D_{{\cal M}_{0},P_{0}}+o_{P}(n^{-1/2}).$
@@
To demonstrate this proposal we consider a statistical model that corresponds to variation independent models of conditional densities. Specifically, let $O=(O_{1},\ldots,O_{J})$ and consider a statistical model ${\cal M}$ that assumes $p(O)=\prod_{j=1}^{J}p_{O(j)\mid Pa(O_{j})}(O(j)\mid Pa(O_{j}))$, but leaves all the conditional densities of $O(j)$, given its parents $Pa(O_{j})$, unspecified, $j=1,\ldots,J$. For example, the likelihood might be factorized according to the ordering $p(O)=\prod_{j=1}^{J}p(O(j)\mid\bar{O}(j-1))$, where $\bar{O}(j-1)=(O(1),\ldots,O(j-1))$, and one might not make any assumptions, or one might assume that for some of the conditional densities $p(O(j)\mid\bar{O}(j-1))$ it is known to depend on $\bar{O}(j-1)$ through a dimension reduction $Pa(O(j))$. This defines then a statistical model ${\cal M}$ for $P_{0}$ only driven by conditional independence assumptions.
Working model defined by estimated dimension reductions: For a given $j$, let $S_{n,j}=S_{j}(P_{n})$ be a data dependent dimension reduction in the sense that $S_{n,j}(Pa(O_{j}))$ is of (much) smaller dimension than $Pa(O_{j})$. Consider the submodel ${\cal M}_{w}$ defined by assuming $p_{O(j)\mid Pa(O_{j})}=p_{O(j)\mid S_{n,j}(Pa(O_{j}))}$, $j=1,\ldots,J$: in other words, this working model assumes that $O(j)$ is independent of $Pa(O_{j})$, given the reduction $S_{n,j}(Pa(O_{j}))$, $j=1,\ldots,J$. Clearly, we have ${\cal M}_{w}\subset{\cal M}$ by making stronger conditional independence assumptions than the ones defining ${\cal M}$.
The Kullback-Leibler projection of $p$ onto ${\cal M}_{w}$ would involve projecting each $p_{j}=p_{O(j)\mid Pa(O)j)}\in{\cal M}_{j}$ onto its smaller working model ${\cal M}_{j,n}$ as follows:
@@
where $L(P)=-\log p$ is the log-likelihood loss. So this projection computes the MLE of $P_{j}$ over the working model under an infinite sample from $P$. Therefore, we have a clear definition of $\Pi_{P}(P\mid{\cal M}_{w})$ and also a clear understanding of how one estimates this projection $\Pi_{P_{0}}(P_{0}\mid{\cal M}_{w})$ with an MLE $\tilde{P}_{n}=\arg\min_{P_{1}\in{\cal M}_{w}}P_{n}L(P_{1})$ over ${\cal M}_{w}$, or, if ${\cal M}_{w}$ is too high dimensional, with a regularized MLE such as an HAL-MLE. If $Pa(O(j))$ is very high dimensional, then an HAL-estimator of the conditional density $p_{0,j}$ is cumbersome, while, if ${\cal M}_{w}$ is indexed by a low dimensional unspecified function, then we can estimate this projection with a powerful HAL-MLE that is also computationally very feasible.